OpenStudy (anonymous):

Solve the following system of equations by using the substitution method. x - y = 3 2x + 2y = 2

OpenStudy (anonymous):

Multiply the first equation by 2 both the sides, tell me what you got?

OpenStudy (anonymous):

@waterineyes 4x+4y=2

OpenStudy (anonymous):

I said multiply 2 to the first equation.. What is your first equation?

OpenStudy (anonymous):

@waterineyes 2x-2y=3

OpenStudy (anonymous):

Why are you not multiplying 2 with right hand side.. I said to multiply both the sides. Now tell me what you get..

OpenStudy (anonymous):

@waterineyes 2x-2y=6

OpenStudy (anonymous):

Yes.. Now add the two equations: 2x - 2y = 6 2x + 2y = 4 Add them..

OpenStudy (anonymous):

@waterineyes 4x+4y=10

OpenStudy (anonymous):

What will be the result when you do this: -2y + 2y = ???

OpenStudy (anonymous):

@waterineyes 0. so y

OpenStudy (anonymous):

i dont understand what you just did

OpenStudy (anonymous):

what about my answer ?? Where it is?

OpenStudy (anonymous):

@muhammad9t5 i didnt see your answer

OpenStudy (anonymous):

@4everjames friend i wrote the answer step by step in 5 minutes.

OpenStudy (anonymous):

@muhammad9t5 what was the answer?

OpenStudy (anonymous):

x=2 and y=-1

OpenStudy (anonymous):

@muhammad9t5 thats what i got. thanks

OpenStudy (anonymous):

@muhammad9t5 Solve the following system of equations by using the elimination method. 2x - 3y = -10 3x + 5y = 4

OpenStudy (anonymous):

@4everjames like my ans please.

OpenStudy (anonymous):

@muhammad9t5 i did

OpenStudy (anonymous):

@4everjames friend do you need ans of 2nd question or explanation?

OpenStudy (anonymous):

@muhammad9t5 both

OpenStudy (anonymous):

friend now it will take time and i am gonna sleep now i wll give expanation tomorrow. please

OpenStudy (anonymous):

@muhammad9t5 ok thanks

OpenStudy (anonymous):

x=-2 and y=2 i hope it helps.